Paul Swanson has an opportunity to acquire a franc
Answer #1

Hi there,

1. Lets first construct the P&L statement :

Sales 510000
Variable expenses:
Cost of ingredients (20%* 510000) 102000
Commissions (14.5%*510000) 73950 175950
Contribution Margin 334050
Selling and administrative expenses
Salaries 91000
Depreciation(396000-39600)/10 35640
Insurance 5600
Utilities 48000
Rent (4800*12) 57600 237840
Net Operating Income $96210

2A) For Simple rate of return:

Rate = NOI/Initial Investment

= 96210/396000 = 24.3%

2B) As the rate of return is higher than the required 22%, he should take the franchise.

3A) To calculate the payback period, we need to calculate the Annual net cash flow

Annual net cash flow = 96210 +35640 = $131850

Payback period = Initial Investment/ Annual net cash flow = 396000/131850 = 3 years

3B) As the payback period is more than 2 years, he will not acquire the franchise.

Kindly rate my answer and give your feedback in the comments section.

Cheers!!